# Improving Online-to-Nonconvex Conversion for Smooth Optimization via Double Optimism

### v2

@@ 1 Introduction
Interestingly, when both the gradient and the Hessian of $F$ are Lipschitz continuous, it becomes possible to outperform standard GD or SGD. In the deterministic setting, [CDHS17] proposed an accelerated gradient-based method that exploits directions of negative curvature, achieving a complexity of $\mathcal{O}(\varepsilon^{-1.75}\log(1/\varepsilon))$. Concurrently, [AABHM17] introduced an approximate variant of the cubic regularization method [NP06], which uses only first-order gradients and access to Hessian-vector products, and achieves a complexity of $\mathcal{O}(\varepsilon^{-1.75}\log(d/\varepsilon))$. More recently, [LL22, LL23] eliminated the logarithmic factor from these bounds by incorporating restarting techniques into accelerated gradient and heavy-ball methods, achieving the state-of-the-art complexity of $\mathcal{O}(\varepsilon^{-1.75})$. However, the methods proposed in [CDHS17, AABHM17] are complex and rely on multiple subroutines to exploit negative curvature, while the restarting techniques in [LL22, LL23] involve lengthy derivations, and their extension to the stochastic setting remains unclear. In the stochastic setting, the best known result is due to [CM20], who proposed normalized SGD with momentum and extrapolation, achieving $\mathcal{O}(\varepsilon^{-2}+\sigma^{2}\varepsilon^{-3.5})$. While this method attains the optimal rate of $\mathcal{O}(\sigma^{2}\varepsilon^{-3.5})$ for the stochastic setting [ACDFSS20], it fails to achieve the accelerated rate of $\mathcal{O}(\varepsilon^{-1.75})$ for the deterministic setting. Notably, to the best of our knowledge, no single unified algorithm simultaneously attains the best-known guarantees in both settings.
@@ 1.1 Additional related work
Apart from the smooth nonconvex optimization considered in this paper, the original O2NC conversion was proposed by [CMO23] also addresses nonsmooth nonconvex optimization, yielding an optimal stochastic algorithm. The framework has been extended and refined in several follow-up works [ZC24, AMC24], and it has been shown that various practical algorithms, such as SGD with momentum and Adam, can be interpreted within the O2NC framework [ZC24, AZKD24, AC24, AMC24].
@@ 3.1 Online optimistic gradient method
To motivate the update rule introduced above, observe that a natural approach would be to base the update of $\bm{\Delta}_{n}$ on the current gradient ${\mathbf{g}}_{n}$, but this is not feasible because ${\mathbf{g}}_{n}$ is revealed only after $\bm{\Delta}_{n}$ has already been selected. To address this challenge, we leverage online optimistic learning techniques from [RS13, JGS20], introducing the first layer of optimism in our doubly optimistic algorithm. These methods rely on a hint vector ${\mathbf{h}}_{n}$ that approximates ${\mathbf{g}}_{n}$ as accurately as possible. The intuition of this approach assumes that the deviation between the true gradient and the hint remains relatively stable across iterations, i.e., ${\mathbf{g}}_{n+1}-{\mathbf{h}}_{n+1}\approx{\mathbf{g}}_{n}-{\mathbf{h}}_{n}$. Accordingly, one can estimate ${\mathbf{g}}_{n+1}\approx{\mathbf{h}}_{n+1}+{\mathbf{g}}_{n}-{\mathbf{h}}_{n}$, and the update combines an optimistic prediction using ${\mathbf{h}}_{n+1}$ with a correction based on the observed error ${\mathbf{g}}_{n}-{\mathbf{h}}_{n}$.
@@
Consider executing the update described in (5) with a constant step size $\eta$. Then the $K$-shifting regret can be bounded by $\mathrm{Reg}_{T}({\mathbf{u}}^{1},\dots,{\mathbf{u}}^{K})\leq\frac{4KD^{2}}{\eta}+\frac{5\eta}{2}\sum_{n=1}^{KT}\|{\mathbf{g}}_{n}-{\mathbf{h}}_{n}\|^{2}-\frac{1}{4\eta}\sum_{n=2}^{KT}\|\bm{\Delta}_{n}-\bm{\Delta}_{n-1}\|^{2}$.
@@ 3.2 Hint construction
For ease of exposition, we first assume access to a deterministic gradient oracle; we will return to the stochastic setting at the end of this section. As explained in Section 3.1, our primary objective in hint construction is to estimate ${\mathbf{g}}_{n}=\nabla F({\mathbf{w}}_{n})$ with maximum accuracy using only information available up to ${\mathbf{x}}_{n-1}$, as illustrated in Figure 1. Leveraging the function’s smoothness properties, an intuitive first approximation is to reuse the previous gradient estimate, assuming ${\mathbf{g}}_{n}\approx{\mathbf{g}}_{n-1}$—effectively positing that the gradient at ${\mathbf{w}}_{n}$ approximates the gradient at ${\mathbf{w}}_{n-1}$. However, this straightforward approximation fails to exploit the negative term $\|\bm{\Delta}_{n}-\bm{\Delta}_{n-1}\|^{2}$ in Lemma 3.1. With this approach, $\|{\mathbf{g}}_{n}-{\mathbf{h}}_{n}\|\leq L\|{\mathbf{w}}_{n}-{\mathbf{w}}_{n-1}\|$ becomes proportional to $\|\bm{\Delta}_{n}+\bm{\Delta}_{n-1}\|$, preventing the application of telescoping arguments. Consequently, this method yields an overall gradient complexity of $\mathcal{O}(\varepsilon^{-11/6})$ in the deterministic setting [CMO23], which is worse than the best-known bound of $\mathcal{O}(\varepsilon^{-7/4})$.
@@
Lemma 3.2.
Consider executing the update described in (5) with the hint construction defined in (6) (or its stochastic counterpart in the stochastic setting), using step size $\eta\leq\frac{1}{\sqrt{3}L_{1}}$. Then we can bound one episode of the $K$-shifting regret, as $\operatorname*{\mathbb{E}}\left[\sum_{n=(k-1)T+1}^{kT}\langle{\mathbf{g}}_{n},\bm{\Delta}_{n}-{\mathbf{u}}^{k}\rangle\right]\leq\frac{4D^{2}}{\eta}+\frac{9L_{1}^{2}\eta D^{2}}{2}+6\eta T\sigma^{2}.$
@@ 3.3 Complexity analysis
Based on Proposition 2.1, achieving the final convergence rate requires bounding the $K$-shifting regret. As a corollary of Lemma 3.2, we get $\operatorname*{\mathbb{E}}\left[\mathrm{Reg}_{T}({\mathbf{u}}^{1},\dots,{\mathbf{u}}^{K})\right]\leq\frac{4KD^{2}}{\eta}+\frac{9L_{1}^{2}\eta KD^{2}}{2}+6\eta KT\sigma^{2}$. Combining this regret bound with Proposition 2.1, we obtain $\operatorname*{\mathbb{E}}\left[\frac{1}{K}\sum_{k=1}^{K}\|\nabla F(\bar{{\mathbf{w}}}^{k})\|\right]\leq\frac{F({\mathbf{x}}_{0})-F^{*}}{DKT}+\frac{4D}{T\eta}+\frac{9L_{1}^{2}\eta D}{2T}+\frac{6\eta}{D}\sigma^{2}+\frac{L_{2}}{48}D^{2}+\frac{L_{2}}{2}T^{2}D^{2}+\frac{\sigma}{\sqrt{T}}.$
@@
We show that after $M$ iterations, the gradient norm at the best iterate satisfies $\min_{1\leq k\leq K}\allowbreak\operatorname*{\mathbb{E}}\left[\|\nabla F(\bar{{\mathbf{w}}}^{k})\|\right]=\mathcal{O}\left(\frac{\sigma^{4/7}}{M^{2/7}}+\frac{1}{M^{4/7}}\right)$, which yields a total gradient complexity of $\mathcal{O}(\sigma^{2}\varepsilon^{-3.5}+\varepsilon^{-1.75})$. This matches the lower bound $\Omega(\sigma^{2}\varepsilon^{-3.5})$ in the stochastic regime [ACDFSW23] and recovers the best known rate $\mathcal{O}(\varepsilon^{-1.75})$ in the deterministic setting, smoothly interpolating between these two regimes.
@@ 4 Adaptive step size scheme
Consider executing the update described in (5), using ${\mathbf{h}}_{n}$ as defined in  (6) (or its stochastic counterpart in the stochastic setting), and the adaptive step size defined in  (8). Then we can bound the first episode of the $K$-shifting regret as $\sum_{n=1}^{T}\operatorname*{\mathbb{E}}\left[\langle{\mathbf{g}}_{n},\bm{\Delta}_{n}-{\mathbf{u}}^{1}\rangle\right]\leq 8\left(\frac{3}{\gamma}+\gamma\right)D\sqrt{T}\sigma+16\left(\frac{3}{\gamma}+\gamma\right)^{\frac{3}{2}}\hat{L}_{1}\gamma^{\frac{1}{2}}D^{2},$ where $\hat{L}_{1}=\max_{1\leq n\leq T}\frac{\|{\mathbf{g}}_{n}-{\mathbf{h}}_{n}\|}{\|{\mathbf{w}}_{n}-{\mathbf{z}}_{n-1}\|}$.
@@ Acknowledgments
This work was supported in part by the NSF CAREER Award CCF-2338846 and the NSF AI Institute for Foundations of Machine Learning (IFML).
@@ Appendix A More details on online-to-nonconvex conversion and related proofs
This bound follows from applying Assumption 2.3, which allows us to establish $\mathbb{E}\bigg[\bigg\|\frac{1}{T}\sum_{n=(k-1)T+1}^{kT}\newline (\nabla F({\mathbf{w}}_{n})-{\mathbf{g}}_{n})\bigg\|\bigg]\leq\frac{\sigma}{\sqrt{T}}$.
@@ Appendix B Proofs for Online Doubly Optimistic Gradient with constant step size
The last inequality follows from Assumption 2.3 that allows us to bound the variance of the first two terms, and the last term comes from leveraging Assumption 2.1, $\big\|\nabla F({\mathbf{x}}_{n-1}+\frac{1}{2}\bm{\Delta}_{n})\newline -\nabla F({\mathbf{x}}_{n-1}+\frac{1}{2}\bm{\Delta}_{n-1})\big\|^{2}\leq L_{1}^{2}\left\|\frac{1}{2}\bm{\Delta}_{n}-\frac{1}{2}\bm{\Delta}_{n-1}\right\|^{2}$. ∎
@@
Note that the bound involves four terms dependent on $T$. If we set $T=\min\bigg(\max\big(\big\lceil\big(\frac{20\sigma}{L_{2}D^{2}}\big)^{\frac{2}{5}}\big\rceil,\allowbreak\big\lceil\big(\frac{10L_{1}}{L_{2}D}\big)^{\frac{1}{3}}\big\rceil\big),\frac{M}{2}\bigg)$ and $K=\lfloor\frac{M}{T}\rfloor$, we can bound the first three terms of the inequality (20) as follows:
@@
  • $2\frac{F({\mathbf{x}}_{0})-F^{*}}{DM}\leq 2\frac{7.5^{\frac{3}{7}}(F({\mathbf{x}}_{0})-F^{*})^{\frac{4}{7}}}{M^{\frac{4}{7}}}+2\frac{16.5^{\frac{5}{7}}(F({\mathbf{x}}_{0})-F^{*})^{\frac{2}{7}}}{M^{\frac{2}{7}}}L_{2}^{\frac{1}{7}}\sigma^{\frac{4}{7}}\leq\frac{5(F({\mathbf{x}}_{0})-F^{*})^{\frac{4}{7}}}{M^{\frac{4}{7}}}L_{1}^{\frac{2}{7}}L_{2}^{\frac{1}{7}}\newline +\frac{15(F({\mathbf{x}}_{0})-F^{*})^{\frac{2}{7}}}{M^{\frac{2}{7}}}L_{2}^{\frac{1}{7}}\sigma^{\frac{4}{7}}$;

### v3

@@ 1 Introduction
Interestingly, when both the gradient and the Hessian of $F$ are Lipschitz continuous, it becomes possible to outperform standard GD or SGD. In the deterministic setting, [CDHS17] proposed an accelerated gradient-based method that exploits directions of negative curvature, achieving a complexity of $\mathcal{O}(\varepsilon^{-1.75}\log(1/\varepsilon))$. Concurrently, [AABHM17] introduced an approximate variant of the cubic regularization method [NP06], which uses only first-order gradients and access to Hessian-vector products, and achieves a complexity of $\mathcal{O}(\varepsilon^{-1.75}\log(d/\varepsilon))$. More recently, [LL22, LL23] eliminated the logarithmic factor from these bounds by incorporating restarting techniques into accelerated gradient and heavy-ball methods, achieving the state-of-the-art complexity of $\mathcal{O}(\varepsilon^{-1.75})$. However, the methods proposed in [CDHS17, AABHM17] are complex and rely on multiple subroutines to exploit negative curvature, while the restarting techniques in [LL22, LL23] involve lengthy derivations, and their extension to the stochastic setting remains unclear. In the stochastic setting, similar improvements over SGD have been achieved using different algorithmic techniques [All18, All18a, TSJRJ18, FLZ19]. To our knowledge, the best known result is due to [CM20], who proposed normalized SGD with momentum and extrapolation, achieving $\mathcal{O}(\varepsilon^{-2}+\sigma^{2}\varepsilon^{-3.5})$. While this method attains the optimal rate of $\mathcal{O}(\sigma^{2}\varepsilon^{-3.5})$ for the stochastic setting [ACDFSS20], it fails to achieve the accelerated rate of $\mathcal{O}(\varepsilon^{-1.75})$ for the deterministic setting. Notably, to the best of our knowledge, no single unified algorithm simultaneously attains the best-known guarantees in both settings.
@@ 1.1 Additional related work
Apart from the smooth nonconvex optimization considered in this paper, the original O2NC conversion by [CMO23] also addressed the complexity of nonsmooth nonconvex optimization first considered in [ZLJSJ20], yielding an optimal stochastic algorithm. The framework has been extended and refined in several follow-up works [ZC24, AMC24], and it has been shown that various practical algorithms, such as SGD with momentum and Adam, can be interpreted within the O2NC framework [ZC24, AZKD24, AC24, AMC24].
@@ 3.1 Online optimistic gradient method
To motivate the update rule introduced above, observe that a natural approach would be to base the update of $\bm{\Delta}_{n+1}$ on the current gradient ${\mathbf{g}}_{n+1}$, but this is not feasible because ${\mathbf{g}}_{n+1}$ is revealed only after $\bm{\Delta}_{n+1}$ has already been selected. To address this challenge, we leverage online optimistic learning techniques from [RS13, JGS20], introducing the first layer of optimism in our doubly optimistic algorithm. These methods rely on a hint vector ${\mathbf{h}}_{n}$ that approximates ${\mathbf{g}}_{n}$ as accurately as possible. The intuition of this approach assumes that the deviation between the true gradient and the hint remains relatively stable across iterations, i.e., ${\mathbf{g}}_{n+1}-{\mathbf{h}}_{n+1}\approx{\mathbf{g}}_{n}-{\mathbf{h}}_{n}$. Accordingly, one can estimate ${\mathbf{g}}_{n+1}\approx{\mathbf{h}}_{n+1}+{\mathbf{g}}_{n}-{\mathbf{h}}_{n}$, and the update combines an optimistic prediction using ${\mathbf{h}}_{n+1}$ with a correction based on the observed error ${\mathbf{g}}_{n}-{\mathbf{h}}_{n}$.
@@
Consider executing the update described in (5) with a constant step size $\eta$. Then the $K$-shifting regret can be bounded by
$\mathrm{Reg}_{T}({\mathbf{u}}^{1},\dots,{\mathbf{u}}^{K})\leq\frac{4KD^{2}}{\eta}+\frac{5\eta}{2}\sum_{n=1}^{KT}\|{\mathbf{g}}_{n}-{\mathbf{h}}_{n}\|^{2}-\frac{1}{4\eta}\sum_{n=2}^{KT}\|\bm{\Delta}_{n}-\bm{\Delta}_{n-1}\|^{2}$
@@ 3.2 Hint construction
For ease of exposition, we first assume access to a deterministic gradient oracle; we will return to the stochastic setting at the end of this section. As explained in Section 3.1, our primary objective in hint construction is to estimate ${\mathbf{g}}_{n}=\nabla F({\mathbf{w}}_{n})$ with maximum accuracy using only information available up to ${\mathbf{x}}_{n-1}$, as illustrated in Figure 1. Leveraging the function’s smoothness properties, an intuitive first approximation is to reuse the previous gradient estimate, assuming ${\mathbf{g}}_{n}\approx{\mathbf{g}}_{n-1}$—effectively positing that the gradient at ${\mathbf{w}}_{n}$ approximates the gradient at ${\mathbf{w}}_{n-1}$. This leads to the following update rule, which corresponds to the standard optimistic online gradient descent
$\bm{\Delta}_{n+1}=\Pi_{\|\bm{\Delta}\|\leq D}\left(\bm{\Delta}_{n}-\eta(2{\mathbf{g}}_{n}-{\mathbf{g}}_{n-1})\right).$
However, this straightforward approximation fails to exploit the negative term $\|\bm{\Delta}_{n}-\bm{\Delta}_{n-1}\|^{2}$ in Lemma 3.1. With this approach, $\|{\mathbf{g}}_{n}-{\mathbf{h}}_{n}\|\leq L\|{\mathbf{w}}_{n}-{\mathbf{w}}_{n-1}\|$ becomes proportional to $\|\bm{\Delta}_{n}+\bm{\Delta}_{n-1}\|$, preventing the application of telescoping arguments. Consequently, this method yields an overall gradient complexity of $\mathcal{O}(\varepsilon^{-11/6})$ in the deterministic setting [CMO23], which is worse than the best-known bound of $\mathcal{O}(\varepsilon^{-7/4})$.
@@
Lemma 3.2.
Consider executing the update described in (5) with the hint construction defined in (6) (or its stochastic counterpart in the stochastic setting), using step size $\eta\leq\frac{1}{\sqrt{3}L_{1}}$. Then we can bound one episode of the $K$-shifting regret, as
$\operatorname*{\mathbb{E}}\left[\sum_{n=(k-1)T+1}^{kT}\langle{\mathbf{g}}_{n},\bm{\Delta}_{n}-{\mathbf{u}}^{k}\rangle\right]\leq\frac{4D^{2}}{\eta}+\frac{9L_{1}^{2}\eta D^{2}}{2}+6\eta T\sigma^{2}.$
@@ 3.3 Complexity analysis
Based on Proposition 2.1, achieving the final convergence rate requires bounding the $K$-shifting regret. As a corollary of Lemma 3.2, we get $\operatorname*{\mathbb{E}}\left[\mathrm{Reg}_{T}({\mathbf{u}}^{1},\dots,{\mathbf{u}}^{K})\right]\leq\frac{4KD^{2}}{\eta}+\frac{9L_{1}^{2}\eta KD^{2}}{2}+6\eta KT\sigma^{2}$. Combining this regret bound with Proposition 2.1, we obtain
$\operatorname*{\mathbb{E}}\left[\frac{1}{K}\sum_{k=1}^{K}\|\nabla F(\bar{{\mathbf{w}}}^{k})\|\right]\leq\frac{F({\mathbf{x}}_{0})-F^{*}}{DKT}+\frac{4D}{T\eta}+\frac{9L_{1}^{2}\eta D}{2T}+\frac{6\eta}{D}\sigma^{2}+\frac{L_{2}}{48}D^{2}+\frac{L_{2}}{2}T^{2}D^{2}+\frac{\sigma}{\sqrt{T}}.$
@@
After $M$ iterations, the best iterate satisfies $\min_{1\leq k\leq K}\operatorname*{\mathbb{E}}\left[\|\nabla F(\bar{{\mathbf{w}}}^{k})\|\right]=\mathcal{O}\left(\frac{\sigma^{4/7}}{M^{2/7}}+\frac{1}{M^{4/7}}\right)$, which yields a total gradient complexity of $\mathcal{O}(\sigma^{2}\varepsilon^{-3.5}+\varepsilon^{-1.75})$. This matches the lower bound $\Omega(\sigma^{2}\varepsilon^{-3.5})$ in the stochastic regime [ACDFSW23] and recovers the best known rate $\mathcal{O}(\varepsilon^{-1.75})$ in the deterministic setting, smoothly interpolating between these two regimes.
Remark 3.1.
While our paper focuses on the smooth setting, Algorithm 1 also recovers the optimal convergence rate in the nonsmooth setting with properly chosen parameters. In this case, we assume that $\mathbb{E}\!\left[\|\nabla f({\mathbf{x}};\xi)\|^{2}\right]\leq G^{2}$ for all ${\mathbf{x}}\in\mathbb{R}^{d}$ and follow the stationary notation of [CMO23]. Under this assumption, Algorithm 1 achieves a complexity of $\mathcal{O}(\varepsilon^{-3}\delta^{-1})$ for finding $(\delta,\varepsilon)$-stationary points. This guarantee is obtained by adapting our analysis to the nonsmooth regime. Specifically, we replace Proposition 2.1 with Theorem 8 of [CMO23], which serves as its nonsmooth analogue, and in the regret bound of Lemma 3.1 we discard the negative term and bound $\|{\mathbf{g}}_{n}-{\mathbf{h}}_{n}\|^{2}$ directly using the bounded second-moment assumption. The remaining steps follow the same structure as the analysis in [CMO23].
@@ 4 Adaptive step size scheme
Consider executing the update described in (5), using ${\mathbf{h}}_{n}$ as defined in  (6) (or its stochastic counterpart in the stochastic setting), and the adaptive step size defined in  (8). Then we can bound the first episode of the $K$-shifting regret as
$\sum_{n=1}^{T}\operatorname*{\mathbb{E}}\left[\langle{\mathbf{g}}_{n},\bm{\Delta}_{n}-{\mathbf{u}}^{1}\rangle\right]\leq 8\left(\frac{3}{\gamma}+\gamma\right)D\sqrt{T}\sigma+16\left(\frac{3}{\gamma}+\gamma\right)^{\frac{3}{2}}\hat{L}_{1}\gamma^{\frac{1}{2}}D^{2},$
where $\hat{L}_{1}=\max_{1\leq n\leq T}\frac{\|{\mathbf{g}}_{n}-{\mathbf{h}}_{n}\|}{\|{\mathbf{w}}_{n}-{\mathbf{z}}_{n-1}\|}$.
@@ Acknowledgments
This work was supported in part by the NSF CAREER Award CCF-2338846, the NSF AI Institute for Foundations of Machine Learning (IFML), and the NSF AI Institute for Future Edge Networks and Distributed Intelligence (AI-EDGE).
@@ Appendix A More details on online-to-nonconvex conversion and related proofs
This bound follows from applying Assumption 2.3, which allows us to establish
$\mathbb{E}\left[\bigg\|\frac{1}{T}\sum_{n=(k-1)T+1}^{kT}\nabla F({\mathbf{w}}_{n})-{\mathbf{g}}_{n}\bigg\|\right]\leq\frac{\sigma}{\sqrt{T}}$.
@@ Appendix B Proofs for Online Doubly Optimistic Gradient with constant step size
The last inequality follows from Assumption 2.3 that allows us to bound the variance of the first two terms, and the last term comes from leveraging Assumption 2.1,
$\left\|\nabla F({\mathbf{x}}_{n-1}+\frac{1}{2}\bm{\Delta}_{n})-\nabla F({\mathbf{x}}_{n-1}+\frac{1}{2}\bm{\Delta}_{n-1})\right\|^{2}\leq L_{1}^{2}\left\|\frac{1}{2}\bm{\Delta}_{n}-\frac{1}{2}\bm{\Delta}_{n-1}\right\|^{2}$
@@
Note that the bound involves four terms dependent on $T$. If we set
$T=\min\left(\max\left(\left\lceil\left(\frac{20\sigma}{L_{2}D^{2}}\right)^{\frac{2}{5}}\right\rceil,\left\lceil\left(\frac{10L_{1}}{L_{2}D}\right)^{\frac{1}{3}}\right\rceil\right),\frac{M}{2}\right),$
and $K=\lfloor\frac{M}{T}\rfloor$, we can bound the first three terms of the inequality (20) as follows:
@@
  • $2\frac{F({\mathbf{x}}_{0})-F^{*}}{DM}\leq 2\frac{7.5^{\frac{3}{7}}(F({\mathbf{x}}_{0})-F^{*})^{\frac{4}{7}}}{M^{\frac{4}{7}}}+2\frac{16.5^{\frac{5}{7}}(F({\mathbf{x}}_{0})-F^{*})^{\frac{2}{7}}}{M^{\frac{2}{7}}}L_{2}^{\frac{1}{7}}\sigma^{\frac{4}{7}}\leq\frac{5(F({\mathbf{x}}_{0})-F^{*})^{\frac{4}{7}}}{M^{\frac{4}{7}}}L_{1}^{\frac{2}{7}}L_{2}^{\frac{1}{7}}+\frac{15(F({\mathbf{x}}_{0})-F^{*})^{\frac{2}{7}}}{M^{\frac{2}{7}}}L_{2}^{\frac{1}{7}}\sigma^{\frac{4}{7}}$;